(Data Analysis Question) Explain supervised machine learning using some examples of data and at least two, two-dimensional supervised machine learning methods. Describe in writing any figures that you normally would draw.
Answer #1

Sion VIS om Supervised machine learning:- Supervised machine learning is one type of humun Suprovision. The main goal in SupeExamples In machine learning, classification is a Supervised learning approach in cohich the computer program leams from theoutlook Nowo, by using cars algorithe, we will solve this problem. Ginile)» 1-603-607)-(3? = 0.643 T P N Gini Index (31) Sann- T N Shot 4 2 P @ 1 3 1 Gini 0.5 0.5 0.315 Scool 4 Gin? (hot) = 1-(2)2-(3) = 0.5 Gani(tild) = 1 -(1)-(1) - 0.5 Go? (cool) =Gini churridity) = 5x0.48 +EXO32 -0.24 +0.16 = 0.40 wird ł Sstrong Sweak 3 7 P N a 5 2 Gini 0.445 0.409 Gini (strong) - 1-(1)c 1 Temp Snot I Shild Scool N 2 T a I 1 P 0 o . Gini 0 o o , 0 0 tu - Ginic Grnictemp) = 0 pod O . . Minimum Minimum , Humidioutlook) - Raindo Sunny brerast (yes (Temp bot / No Rainy N :) T SRainy 3 a Gini 0.445 Gini se 0 S T strong 1 Sweak a D , Ń 0(outlook) . Rainy Overcast - Sunny / Temp yes) weak ong Mild hot no yesand Examplet using linear Regression ao ave an Economist for the Co-operative. you gasher the follocoing data: Fertilizer(lb)bo = 9 - 8,1 = 6 – (0-65)() -0.80 Y = 0.8 +0.6534 10+ G=0.8 +0.65%, Yeld(123) 67 0 ut .. . Fertilizer (lb) a-dimensional Machbo = 9 - 8,1 = 6 – (0-65)() -0.80 Y = 0.8 +0.6534 10+ G=0.8 +0.65%, Yeld(123) 67 0 ut .. . Fertilizer (lb) a-dimensional Machpositive linear Relationship Negative linear Relationship Ely) Ely) Regression line Regression line Ilhi Intercept slope B, iliniquely plentifies the straight line passes through these 2 points ) Support Vector Machines!. this method is also 2-dimensExamples of keinel fonctions feln polynomial kernel with degreed K(x,y) = (xTy tijd Radial basis function Kernel with width o

please rate with up vote thank you